OpenStudy (sasogeek):

It was Helen's first day at school. The teacher suggested that it would be a good idea for each child to meet every other child in the class. The teacher said, "When you meet, please shake hands and introduce yourself by name." If there were 10 children in the class, how many total handshakes were there?

OpenStudy (anonymous):

Don't you want to challenge me? :P

OpenStudy (ash2326):

This is just like selecting any two people out of 10 = 10C2 or 45

OpenStudy (sasogeek):

LOL soon ffm

OpenStudy (anonymous):

\(\Huge \frac{n\times (n-1)}{2} \)

OpenStudy (anonymous):

Yes, 45. It is like 1+2+3+....+9.

OpenStudy (anonymous):

\(\Huge\sum \limits_{i=1}^{n-1} i \)